From GlueXWiki
Jump to: navigation, search

July 24, 2014 FDC+CDC meeting


  1. DAQ (Beni)
  2. CDC status (Mike, Beni)
  3. FDC status First Track through all packages (Lubomir)
  4. Electronics (Chris, Nick)
  5. Cathode scan (Luke)
  6. Other

Instructions for Bluejeans meeting connection

FDC meeting ID: 290664653


Participants: Sean (NU), Luke, Chris, Mike, Simon, Beni, Fernando, Eugene, and Lubomir (JLab).


- Beni: DAQ doesn't want to crash now, so it's difficult to troubleshoot the problem. Will start big run over the weekend with the new (nonflammable) cable to record the trigger time.

CDC update

- Beni studied the CDC noise once again (all the results are linked above): using a scope on one channel (#4 connector C6) with a stand alone DC power. He made measurements in different conditions: turning on/off LV/HV the power for the crates itself and also ramping the HV. He was looking for is 60Hz noise and 200kHz bursts with tens of MHz within the bursts. The conclusion: the 60Hz comes when powering the south HV supply (connected to the card). The 200kHz bursts come when rumping up the HV and stay when the HV is up. Fernando wanted more tests with more cards and HV modules to see if this is a problem with a particular module/HV crate or a common feature of this CAEN system.

- Mike found that for the "problematic" channels (with no signals in the fADC but visible on the scope) one can actually see something (noise or signal) in the fADC but with a huge DAQ offset of 0xd000. It is not clear yet what is the quality of the signals.

FDC update

- The first track going through all the 24 chambers is shown above. It was obtained with Beni's horizontal muon trigger. So far within 4 hours run we found only two such events. There are few more with muons partially crossing the FDC system. Still, this one event help a lot to fix the geometry, the translation table and some bugs in the code. Important is that we see signals from all the cells and all the planes within a cell, upper strips, down strips, and a wire, crossing at the same point, meaning we have 100% efficiency. The space points in the plot are the avalanche positions obtained from the cathodes only and fitted with a 3d line. Next step is to use the drift time information to find the actual hit positions.


- Chris and Nick worked on the CDC testing the "problematic" channels. Most likely the problem is in the 3M connector (but could be somewhere else), so Chris is going to exercise to see if one can open and fix such connector.

- Fernando: Cody is ready with the pulser firmware for fADC125. After some iterations with Bryan, it will be ready for use.

Cathode scan

- Luke finished the first set of measurements: for a fixed y (along the strips at about 20cm from center) he scanned both sides of all the strips. The data can be explained by using a regular pitch of 5.012 mm without any gaps or change of the pitch values in the different foils. There are still some "jumps" off the expected values in the splicing regions but of the order of 50 microns. Eugene: how good is the repeatability? Luke: we are counting now the motor revolutions (encoder not connected yet), but if you go in one direction, as the measurements were done, it is accurate to 20 microns. Luke will continue with the scan at other y-values, will make distributions of strip widths and other parameters.


- Scot replaced the faulty change-over unit for the CO2 on Wednesday. The chiller failed on Friday ( "alarm 12: return low flow flt"), reset didn't help, had to unplug/plug the power. Wesley Moore was working before that using the serial port for the controls, but not when it failed.